# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

### v2

@@ V Conclusions
This paper proposed a fRIS-aided ISAC system to achieve multi-user communications and multi-target sensing by exploiting the beamforming and the fRIS element position optimization. We formulated a joint communication symbol estimation MSE and sensing beampattern mismatch minimization problem subject to transmit power budget, constant-modulus constraint and the element position requirements of the fRIS. An AM-based algorithm was developed, where each subproblem utilized the techniques of derivatives, SDR, ALM and MM optimization. Simulation results validated that the fRIS can significantly reduce the sensing beampattern mismatch and communication symbol estimation error in the ISAC systems, as compared to the conventional RIS.
@@ Appendix A Derivation of the function $\widetilde{f}_{0}$ in (III-E)
To simplify the objective function (34a), we first expand the two square norms, and it is given as
@@
It is noteworthy that $\mathbf{s}_{r}^{\mathrm{H}}\mathbf{s}_{r}$ and $\mathbf{s}_{c}^{\mathrm{H}}\mathbf{s}_{c}$ are constants, which have been omitted.
Observing the first term $\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$, it is actually unrelated to the variable $\mathbf{p}$ due to the following reason. Given the property of $\boldsymbol{\Theta}$ that $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ and the expression of $\mathbf{G}$ in (10), this term can be written as $\zeta_{G}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$. Since $\mathbf{a}_{r}(\mathbf{p})$ has the structure of array manifold of (8), $\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})=N$ always holds regardless of element positions. As a result, this term eventually becomes $\zeta_{G}N\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$, which is unrelated to $\mathbf{p}$ and ignorable.
For the second term $-2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}$, the expression of $\mathbf{G}$ can also be substituted, yielding $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\}$. Leveraging the cyclic property of trace, it can be re-expressed as $-2\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, where $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$.
Focusing on the third term, it can be rewritten as
@@
In operation (a), the expression of $\mathbf{G}$ in (10) and $\mathbf{H}_{rc}$ in (11) are substituted, while $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$ and $c_{0}=\omega^{2}\zeta_{G}c_{x}$ incorporate the variable-independent terms for simplicity. In operation (b), it is derived based on the matrix structure in (14a) and (14b). The operation (c) leverages the diagonal property of $\boldsymbol{\Theta}$, where we denote $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$.
As for the last term in (A), it can be transformed as follows:
@@
where (a) is the substitution operation, while (b) utilizes the cyclic property of trace and the diagonal property of $\boldsymbol{\Theta}$. We also denote $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
@@ Appendix B Derivation of the function $\widetilde{f}_{1}$ in (III-E)
For the function (III-E), it can be further simplified as follows if one considers the position of $n$-th element position $\mathbf{p}_{n}$ while fixing others. First, examining the first term $-2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, it is the sum of $N$ number, which can be expressed as
@@
[…]
In particular, $[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}$ is $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\rho_{n}=|[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}|$ and $\xi_{n}=\angle([\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n})$. Additionally, $\nu_{n}=2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is a variable-independent value. The symbol $d_{r,n}(\mathbf{p}_{n})$ is determined by $\mathbf{p}_{n}$, which is given as $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})+p_{y,n}\sin(\psi_{r})$. Moreover, in the operation (a), only the term related to $\mathbf{p}_{n}$ is kept since other element positions are considered constant.
For the second term, one can rewrite it in element-wise manner, i.e.,
@@
where $d_{c,k,n}(\mathbf{p}_{n})=p_{x,n}\sin(\phi_{c,k})\cos(\psi_{c,k})+p_{y,n}\sin(\psi_{c,k})$. Apparently, when $i=j$, the complex exponential terms are mutually canceled out and become constants unrelated to $\mathbf{p}$. On the other hand, it can be found that the terms of $i\neq j$ appear in conjugate pairs since $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}$ is a Hermitian matrix, Owing to this characteristics, when considering $n$-th element (i.e. $i=n\text{ or }j=n$) and ignoring the constant terms of $i=j$, (53c) can be reduced to
@@
Notably, when fixing the positions of other elements, the term $[\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}]_{i,n}\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}$ is unrelated to $\mathbf{p}_{n}$, which can be viewed constant as $\widetilde{c}_{i,n,k}$. We denote the phase of $\widetilde{c}_{i,n,k}$ as $\widetilde{\xi}_{i,n,k}=\angle\widetilde{c}_{i,n,k}$ while $\widetilde{c}_{i,n,k}$ has an amplitude of $1$. As a result, (B) can further be written as
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\!\zeta_{k}\!\!\!\!\sum_{i=1,i\neq n}^{N}\!\!\!\!2\mathfrak{R}\left\{\widetilde{c}_{i,n,k}e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}$ (55a)
$\longrightarrow$ $\sum_{k=1}^{K}\!\sum_{i=1,i\neq n}^{N}\!\!\widetilde{\nu}_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (55b)
where we denote $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})=d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\nu}_{k}=2(1\!-\!\alpha)c_{0}\zeta_{k}$ for simplicity.
@@
$\longrightarrow$ $\!\sum_{n=1}^{N}\sum_{k=1}^{K}\!\!-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n}))}\right\}$ (56b)
$\longrightarrow$ $\sum_{k=1}^{K}-2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n}))}\right\}$ (56c)
$\longrightarrow$ $\sum_{k=1}^{K}\!-\overline{\nu}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (56d)
where $s_{n,k}=[\mathbf{S}]_{n,k}$ with the amplitude of $\overline{\rho}_{n,k}=|s_{n,k}|$ and the phase of $\overline{\xi}_{n,k}=\angle s_{n,k}$. We also denote $\overline{\nu}_{n,k}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}\overline{\rho}_{n,k}$ for brevity.
@@ Appendix C Gradient and Hessian matrix Derivation of $\widetilde{f}_{1}$
For the first sub-function $\widetilde{f}_{1,1}$, the first-order derivative with respect to $p_{x,n}$ and $p_{y,n}$ is the derivative of a composite function, where the outer function is a cosine function while the interior function is linear. As a result, the can be handled by the chain rule of derivative. First, we derive the the first-order derivatives of $\widetilde{f}_{1,1}$:
@@
$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{x,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,x}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57a)
$\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{y,n}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\kappa_{r,y}\sin\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (57b)
@@
where $\kappa_{r,x}=\sin(\phi_{r})\cos(\psi_{r})$ and $\kappa_{r,y}=\sin(\psi_{r})$. Based on (57), the second-order derivatives can be derived as
@@
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58a)
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}\partial p_{y,n}}\!$ $=\!\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}\partial p_{x,n}}$
$=\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,x}\kappa_{r,y}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right),$ (58b)
$\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}^{2}}\!$ $=\!\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\kappa_{r,y}^{2}\cos\!\left(\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}\!\right).$ (58c)
For the second sub-function $\widetilde{f}_{1,2}$, the first-order derivatives and the second-order derivatives can also be derived as
@@
$\!\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{x,n}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59a)
@@
$\!\frac{\partial\widetilde{f}_{1,2}}{\partial p_{y,n}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{y,k}\sin\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right),$ (59b)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}^{2}}\!=\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59c)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}\partial p_{y,n}}\!=\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}\partial p_{x,n}}$
@@
$\!=\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{x,k}\Delta\widetilde{\kappa}_{y,k}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59d)
$\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}^{2}}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\widetilde{\kappa}_{y,k}^{2}\cos\!\left(\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\!\right)\!,$ (59e)
where $\Delta\widetilde{\kappa}_{x,k}=\kappa_{r,x}-\kappa_{c,x,k}$ and $\kappa_{c,x,k}=\sin(\phi_{c,k})\cos(\psi_{c,k})$, while $\Delta\widetilde{\kappa}_{y,k}=\kappa_{r,y}-\kappa_{c,y,k}$ and $\kappa_{c,y,k}=\sin(\psi_{c,k})$.
Similarly, the first-order and second-order derivatives of $\widetilde{f}_{1,3}$ can be given as
@@
$\!\frac{\partial\widetilde{f}_{1,3}}{\partial p_{x,n}}\!=\!\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\widetilde{\kappa}_{x,k}\sin\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}\right),$ (60a)

### v3

@@ V Conclusions
This paper proposed a fRIS-aided ISAC system to achieve multi-user communication and multi-target sensing, where a joint communication MSE and sensing beampattern mismatch minimization problem was formulated. To address the problem, an AM-based algorithm was developed, which involved various techniques. Simulation results validated that the fRIS significantly reduces the sensing beampattern mismatch and communication BER. Future research may explore new fRIS architectures in ISAC systems, such as active or beyond diagonal fRIS designs. In addition, it is interesting to study the modeling of fRIS under rich-scattering environments.
@@ Appendix A Derivation of the function $\widetilde{f}_{0}$ in (III-E)
To simplify (31a), we first expand the two square norms, which is given by
@@
Here, $\mathbf{s}_{r}^{\mathrm{H}}\mathbf{s}_{r}$ and $\mathbf{s}_{c}^{\mathrm{H}}\mathbf{s}_{c}$ are constants, which have been omitted.
Observing the term $\mathbf{x}^{\mathrm{H}}\mathbf{G}^{\mathrm{H}}\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}$, it is unrelated to $\mathbf{p}$ due to the following reason. Given $\boldsymbol{\Theta}\boldsymbol{\Theta}^{\mathrm{H}}=\mathbf{I}$ and the expression of $\mathbf{G}$ in (9), this term can be written as $\zeta_{G}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$. Since $\mathbf{a}_{r}(\mathbf{p})$ has the structure of array manifold of (7), $\mathbf{a}_{r}^{\mathrm{H}}(\mathbf{p})\mathbf{a}_{r}(\mathbf{p})=N$ always holds regardless of $\mathbf{p}$. Therefore, this term becomes $\zeta_{G}N\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}$, which is unrelated to $\mathbf{p}$ and ignorable.
For the term $2\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{G}\mathbf{x}\}$, the expression of $\mathbf{G}$ can be substituted, yielding $2\sqrt{\zeta_{G}}\mathfrak{R}\{\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\}$. Leveraging the cyclic property of trace, it is re-expressed as $2\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, where $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{r}^{\mathrm{H}}\boldsymbol{\Theta}^{\mathrm{H}}$.
Focusing on the third term, it is rewritten as
@@
In (a), the expression of $\mathbf{G}$ and $\mathbf{H}_{rc}$ are substituted, while $c_{x}=\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{x}^{\mathrm{H}}\mathbf{a}_{t}$ and $c_{0}=\omega^{2}\zeta_{G}c_{x}$ incorporate the variable-independent terms. In (b), it is derived based on the matrix structure in (13a) and (13b). The operation (c) leverages the diagonal property of $\boldsymbol{\Theta}$, where we denote $\widetilde{\mathbf{A}}_{c,k}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{c,k}(\mathbf{p}))$.
For the last term in (A), it is transformed as follows:
@@
where (a) is the substitution operation, while (b) utilizes the cyclic property of trace and the diagonal property of $\boldsymbol{\Theta}$. In addition, we denote $\widetilde{\mathbf{A}}_{r}(\mathbf{p})=\mathrm{Diag}(\mathbf{a}_{r}(\mathbf{p}))$ and $\mathbf{S}=\boldsymbol{\theta}^{*}\mathbf{a}_{t}^{\mathrm{H}}\mathbf{x}\mathbf{s}_{c}^{\mathrm{H}}\boldsymbol{\Sigma}_{rc}^{\mathrm{H}}$.
@@ Appendix B Derivation of the function $\widetilde{f}_{1}$ in (III-E)
For (III-E), it can be simplified if we consider the position optimization of the $n$-th element $\mathbf{p}_{n}$ while fixing others. First, examining the first term $2\alpha\sqrt{\zeta_{G}}\mathfrak{R}\{\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}\mathbf{a}_{r}(\mathbf{p})\}$, it is the sum of $N$ number, which is expressed as
@@
${\longrightarrow}$ $-\nu_{n}\cos\left(\xi_{n}+\frac{2\pi}{\lambda}d_{r,n}(\mathbf{p}_{n})\right).$ (46c)
In particular, $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}$ is the $n$-th entry of $\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}$, while $\rho_{n}=|[\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n}|$ and $\xi_{n}=\angle([\widetilde{\mathbf{a}}_{t}^{\mathrm{H}}]_{n})$. Additionally, $\nu_{n}=2\alpha\sqrt{\zeta_{G}}\rho_{n}$ is a variable-independent value. The symbol $d_{r,n}(\mathbf{p}_{n})$ is determined by $\mathbf{p}_{n}$, i.e., $p_{x,n}\sin(\phi_{r})\cos(\psi_{r})+p_{y,n}\sin(\psi_{r})$. Moreover, in (a), only the term related to $\mathbf{p}_{n}$ is kept since other element positions are considered constant.
For the second term, it can be rewritten in an element-wise manner, i.e.,
@@
where $d_{c,k,n}(\mathbf{p}_{n})=p_{x,n}\sin(\phi_{c,k})\cos(\psi_{c,k})+p_{y,n}\sin(\psi_{c,k})$. When $i=j$, the complex exponential terms are mutually canceled out and become constants unrelated to $\mathbf{p}$. Besides, the terms of $i\neq j$ appear in conjugate pairs since $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}$ is a Hermitian matrix, Owing to this characteristics, when considering the $n$-th element (i.e. $i=n\text{ or }j=n$) and ignoring the constant terms of $i=j$, (47c) is reduced to
@@
Notably, when fixing the positions of other elements, the term $\boldsymbol{\theta}\boldsymbol{\theta}^{\mathrm{H}}]_{i,n}\cdot e^{-\jmath\frac{2\pi}{\lambda}\left(d_{r,i}(\mathbf{p}_{i})-d_{c,k,i}(\mathbf{p}_{i})\right)}$ is unrelated to $\mathbf{p}_{n}$, which can be viewed constant as $\widetilde{c}_{i,n,k}$. We denote the phase of $\widetilde{c}_{i,n,k}$ as $\widetilde{\xi}_{i,n,k}=\angle\widetilde{c}_{i,n,k}$ while $\widetilde{c}_{i,n,k}$ has an amplitude of $1$. As a result, (B) can be written as
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!2(1\!-\!\alpha)c_{0}\sum_{k=1}^{K}\!\zeta_{k}\!\!\!\!\sum_{i=1,i\neq n}^{N}\!\!\!\!\!\mathfrak{R}\left\{\widetilde{c}_{i,n,k}e^{\jmath\frac{2\pi}{\lambda}\left(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})\right)}\right\}$ (49a)
$\longrightarrow$ $\sum_{k=1}^{K}\!\sum_{i=1,i\neq n}^{N}\!\!\widetilde{\nu}\zeta_{k}\cos\left(\widetilde{\xi}_{i,n,k}+\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (49b)
where we denote $\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})=d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n})$ and $\widetilde{\nu}=2(1\!-\!\alpha)c_{0}$ for simplicity.
@@
$\longrightarrow$ $\!\sum_{n=1}^{N}\sum_{k=1}^{K}\!\!-\overline{\nu}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}(d_{r,n}(\mathbf{p}_{n})-d_{c,k,n}(\mathbf{p}_{n}))}\right\}$ (50b)
$\longrightarrow$ $\sum_{k=1}^{K}-\overline{\nu}\mathfrak{R}\left\{s_{n,k}e^{\jmath\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n}))}\right\}$ (50c)
$\longrightarrow$ $\sum_{k=1}^{K}\!-\overline{\nu}~\overline{\rho}_{n,k}\cos\!\left(\overline{\xi}_{n,k}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}(\mathbf{p}_{n})\right),$ (50d)
where $s_{n,k}=[\mathbf{S}]_{n,k}$ with the amplitude of $\overline{\rho}_{n,k}=|s_{n,k}|$ and the phase of $\overline{\xi}_{n,k}=\angle s_{n,k}$. We also denote $\overline{\nu}=2(1\!-\!\alpha)\omega\sqrt{\zeta_{G}}$ for brevity.
@@ Appendix C Gradient and Hessian matrix Derivation of $\widetilde{f}_{1}$
For the sub-function $\widetilde{f}_{1,1}$, the first-order derivative w.r.t. $p_{x,n}$ and $p_{y,n}$ is the derivative of a composite function, where the outer function is a cosine function while the interior function is linear. Therefore, we can adopt the chain rule of derivatives. First, we derive the the first-order derivatives of $\widetilde{f}_{1,1}$:
@@
$\left[\!\frac{\partial\widetilde{f}_{1,1}}{\partial p_{x,n}},\frac{\partial\widetilde{f}_{1,1}}{\partial p_{y,n}}\right]^{\mathrm{T}}\!=\!\nu_{n}\frac{2\pi}{\lambda}\boldsymbol{\kappa}_{r}\sin\!\left(\tau_{n}\right),$ (51a)
@@
where $\boldsymbol{\kappa}_{r}=[\kappa_{r,x},\kappa_{r,x}]^{\mathrm{T}}=[\sin(\phi_{r})\cos(\psi_{r}),\sin(\psi_{r})]^{\mathrm{T}}$ and $\tau_{n}=\xi_{n}\!+\!\frac{2\pi}{\lambda}d_{r,n}$. Based on (51), the second-order derivatives are derived as
@@
$\begin{bmatrix}\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}^{2}}&\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{x,n}\partial p_{y,n}}\\ \frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}\partial p_{x,n}}&\frac{\partial^{2}\widetilde{f}_{1,1}}{\partial p_{y,n}^{2}}\end{bmatrix}=\nu_{n}\frac{4\pi^{2}}{\lambda^{2}}\cos\!\left(\tau_{n}\right)\boldsymbol{\kappa}_{r}\boldsymbol{\kappa}_{r}^{\mathrm{T}}.$ (52)
For the second sub-function $\widetilde{f}_{1,2}$, the first-order derivatives and the second-order derivatives can be derived as
@@
$\!\!\!\left[\frac{\partial\widetilde{f}_{1,2}}{\partial p_{x,n}},\frac{\partial\widetilde{f}_{1,2}}{\partial p_{y,n}}\right]=\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!\!-\widetilde{\nu}_{k}\frac{2\pi}{\lambda}\Delta\boldsymbol{\widetilde{\kappa}}_{k}\sin\!\left(\widetilde{\tau}_{i,k,n}\right)$ (53a)
@@
$\!\!\!\begin{bmatrix}\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}^{2}}&\!\!\!\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{x,n}\partial p_{y,n}}\\ \frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}\partial p_{x,n}}&\!\!\!\!\!\!\frac{\partial^{2}\widetilde{f}_{1,2}}{\partial p_{y,n}^{2}}\end{bmatrix}\!=\!\!\sum_{i\neq n}^{N}\sum_{k=1}^{K}\!-\widetilde{\nu}_{k}\frac{4\pi^{2}}{\lambda^{2}}\Delta\mathbf{K}_{k}\cos\left(\widetilde{\tau}_{i,k,n}\right)\!,$ (53b)
@@
where $\widetilde{\tau}_{i,k,n}=\!\widetilde{\xi}_{i,n}\!+\!\frac{2\pi}{\lambda}\Delta\widetilde{d}_{k,n}$, $\Delta\mathbf{K}_{k}=\Delta\boldsymbol{\widetilde{\kappa}}_{k}\Delta\boldsymbol{\widetilde{\kappa}}_{k}^{\mathrm{T}}$, $\Delta\boldsymbol{\widetilde{\kappa}}_{k}=\boldsymbol{\kappa}_{r}-\boldsymbol{\kappa}_{c,k}$ and $\boldsymbol{\kappa}_{c,k}=[\sin(\phi_{c,k})\cos(\psi_{c,k}),\sin(\psi_{c,k})]^{\mathrm{T}}$.
Similarly, the first-order and second-order derivatives of $\widetilde{f}_{1,3}$ can be given by
@@
$\!\!\!\!\!\left[\frac{\partial\widetilde{f}_{1,3}}{\partial p_{x,n}},\frac{\partial\widetilde{f}_{1,3}}{\partial p_{y,n}}\right]=\sum_{k=1}^{K}\overline{\nu}_{n,k}\frac{2\pi}{\lambda}\Delta\boldsymbol{\widetilde{\kappa}}_{k}\sin\!\left(\overline{\tau}_{n,k}\right)$ (54a)